OpenStudy (anonymous):

Essay; show all work. A simple dartboard has three areas… the main board has a radius of 8 inches, there is a circle with a radius of 6 inches, and the bullseye has a radius of 4 inches. What is the probability of a random dart landing inside the bullseye?

OpenStudy (anonymous):

from what i am reading onling you Pi * 4(2) ________ pi* 8(2) which equals 16/64=1/4 or 25% but how are they doing this?

OpenStudy (anonymous):

is there a formula i cannot find?

OpenStudy (unklerhaukus):

You might have to assume that the dart its the board , the sum of probabilities is 1, i guess the probability of landing inside one area is proportional to its relative area on the whole board

OpenStudy (anonymous):

the answer seems to be correct, but what i dont understand is how does pi *4(2) squared ------------ pi * 8(2) squared = 16/64=1/4 wouldnt you multiply pi to them insteasd of just squaring 4 and 8?

OpenStudy (unklerhaukus):

the area of a circle \[A=\pi r^2\] for a radius of 4 \[A_{Bulls Eye}=\pi(4)^2=16\pi\] for a radius of 8 \[A_{Board}=\pi(8)^2=64\pi\] the ratio of the bulls eye to the total board is \[\frac{A_{Bulls Eye}}{A_{board}}=\frac{16\pi}{64\pi}=\frac{ 1}{ 4}\]

OpenStudy (unklerhaukus):

Are you are those measurements are of the radius? a one in four chance of bulls eye seems too large to me , perhaps the measurements are of the diameter?

OpenStudy (anonymous):

Essay; show all work. A simple dartboard has three areas… the main board has a radius of 8 inches, there is a circle with a radius of 6 inches, and the bullseye has a radius of 4 inches. What is the probability of a random dart landing inside the bullseye? Round to the nearest thousandth. (Hint: Use A=pir2

OpenStudy (unklerhaukus):

ok so it just an easy dart board with a huge bulls eye

OpenStudy (anonymous):

Must be lol but thank you so much for your help!! its so appreciated!